OpenStudy (anonymous):

1/2(16-4a) - 3/4(12+20a) Help!!

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }(16-4a)-\frac{ 3 }{ 4 }(12+20a)\]

OpenStudy (timaashorty):

\[\frac{ 1 }{ 2 }(16-4a) - \frac{ 3 }{ 4 }(12+20a) = \]\[\frac{ 1 }{ 2} (4) - \frac{ 3 }{ 4 } ( \frac{ -3 }{ 5 }) =\] \[a=-\frac{ 1 }{ 17 }\]